# When Personas Override Payoffs: Role Identity Bias in Multi-Agent LLM Decision-Making

### v2

@@ Abstract.
Large language models are increasingly deployed in multi-agent systems for strategic tasks, yet how design choices such as role-based personas and payoff visibility affect their reasoning capabilities remains poorly understood. We investigate whether multi-agent systems function as strategic reasoners capable of payoff optimization or as identity-driven actors that prioritize role alignment over explicit incentives. We use Nash equilibrium achievement in game-theoretic settings as a diagnostic tool to assess reasoning capabilities. We do so through systematic experiments with four Large language model (LLM) architectures (Qwen-7B/32B, Llama-8B, Mistral-7B) across complex environmental decision-making scenarios in four-agent strategic games. We demonstrate that role identity bias fundamentally alters strategic reasoning, i.e. when personas are active, all models achieve near-zero Nash equilibrium rates (0–6.7%) in scenarios where payoff-optimal equilibria exist, despite having complete payoff information. Removing personas and providing explicit payoffs enables Qwen models to achieve high Nash equilibrium rates (65–90%) in these scenarios, demonstrating that both conditions are necessary for strategic reasoning capabilities. Personas systematically bias equilibrium selection toward socially preferred outcomes, with 100% of Nash equilibria being Green Transition (socially preferred) when personas are present, while models completely fail to reach equilibrium in scenarios where Tragedy is optimal. The effect of explicit payoffs depends entirely on persona presence, revealing that design choices about personas and payoff visibility interact in ways that fundamentally determine system behavior. We observe striking model-dependent patterns where Qwen architectures show high sensitivity to both personas and payoff visibility, while Llama and Mistral maintain rigid reasoning patterns across conditions. These findings establish that representational choices are substantive governance decisions that determine whether multi-agent systems function as strategic reasoners or identity-driven actors, with important implications for deployment in strategic settings.
@@ 1. Introduction
Multi-agent systems powered by large language models are increasingly being deployed in policy simulation, economic modeling, negotiation platforms and other settings that require strategic reasoning (Merrill and Srivastava, 2025; Li et al., 2024; Bianchi et al., 2024a). In these applications, agents are frequently assigned role-based personas to represent different stakeholders or perspectives (Merrill and Srivastava, 2025; Xu et al., 2024b). These personas enable more realistic simulations and help systems model complex multi-stakeholder scenarios (Park et al., 2023), particularly in environmental and climate-related decision-making, where heterogeneous stakeholders interact strategically and coordination failures can lead to persistent social inefficiencies (Hardin, 1968; Ostrom, 1990; Barrett, 2003; Reyes-Cruz et al., 2024). However, a critical question remains: when agents with role-based personas face scenarios that demand strategic decision making, do they follow explicit payoffs or their assigned identities?
This question determines whether multi-agent large language model (LLM) systems behave as strategic reasoners or as identity-driven actors who prioritize role alignment over payoff maximization (Herr et al., 2024; Duan et al., 2024b; Neumann et al., 2025). If personas override explicit payoff information, a system designed for strategic analysis might instead follow the persona-consistent behavior that fails to optimize over encoded incentives (Neumann et al., 2025; Herr et al., 2024). This creates a fundamental tension between the representational goals of persona-based systems and their ability to reason strategically when payoff information is available (Bender et al., 2021; Duan et al., 2024b). Understanding how personas interact with payoff information in multi-agent settings is essential to make deployment decisions (Weidinger et al., 2022b; Chan et al., 2024; Casper et al., 2024; Rivera et al., 2024; Chan et al., 2024). However, this gap remains largely unexplored. Existing work has examined personas, payoff visibility, and model differences separately but has not systematically studied how these factors together determine system behavior in strategic settings (Neumann et al., 2025; Duan et al., 2024b; Akata et al., 2025; Huang et al., 2025; Sun et al., 2025; Zhang et al., 2024).
Refer to caption Flow diagram showing four agents with chain-of-thought reasoning making simultaneous action selections A flow diagram illustrating the multi-agent game structure with four agents (Environmentalist, Government, Industrialist, Citizen) engaging in simultaneous action selection. Each agent is represented by a robot figure with a dotted speech bubble above it containing chain-of-thought (CoT) reasoning text that shows how the agent analyzes the scenario and their role. The Environmentalist’s CoT begins with ”From the scenario, it is clear that…”, the Government’s with ”The scenario suggests that…”, the Industrialist’s with ”As an Industrialist, I need to…”, and the Citizen’s with ”Given my role as a citizen, I must…”. Each agent has two binary action choices (shown on arrows): Environmentalist chooses CAMPAIGN or NOCAMPAIGN, Government chooses REGULATE or NOREG, Industrialist chooses POLLUTE or CLEAN, and Citizen chooses SUPPORT_GREEN or BUY_CHEAP. All four arrows converge into a central ”Strategy Profile” box, representing the simultaneous action selection where all agents choose their actions at the same time, and the combination of all four actions forms the final strategy profile that is evaluated for Nash equilibrium.
Figure 1. The diagram shows our multi-agent game structure with chain-of-thought reasoning. Each agent analyzes the scenario and their role identity, then selects one of two binary actions. The arrows from each agent converge to a central Strategy Profile box, representing simultaneous action selection where all four agents choose actions concurrently, and their combined choices form the strategy profile evaluated for Nash equilibrium. The dotted speech bubbles contain each agent’s chain-of-thought reasoning process.
Understanding how role identity and payoff optimization interact is critical in multi-agent strategic games. We study environmental policy scenarios with four agents (Industrialist, Government, Environmental Activist, and Citizen) coordinating to reach Nash equilibria (Nash, 1951), which can yield either Green Transition (socially preferred and environmentally sustainable) or Tragedy of Commons (payoff-optimal but environmentally damaging) outcomes (Hardin, 1968). Through a $\times2$ factorial experiment manipulating persona presence and payoff visibility across four model architectures, we examine whether agents prioritize role-aligned actions or payoff-optimal strategies. Our findings show that personas fundamentally alter strategic reasoning, preventing payoff optimization: when personas are active, all models achieve near-zero Nash equilibrium rates in economic scenarios where Tragedy of Commons is payoff-optimal, despite complete payoff information. Removing personas and providing explicit payoffs enables certain models to achieve high Nash equilibrium in these scenarios, demonstrating both conditions are necessary for strategic reasoning. Personas systematically bias equilibrium selection toward socially preferred outcomes, with Nash equilibria being exclusively Green Transition when personas are present, while models fail to reach equilibrium in scenarios where Tragedy is optimal. The effect of explicit payoffs depends entirely on persona presence, and we observe model-dependent patterns where some architectures show high sensitivity to both personas and payoff visibility, while others maintain rigid reasoning patterns across conditions. This work provides key contributions to understanding LLM behavior in multi-agent strategic settings. We demonstrate that role identity bias completely overrides explicit payoff optimization, preventing Nash equilibrium. Additionally, personas and payoff visibility interact in ways that fundamentally determine system behavior, meaning design choices about representation formats are substantive governance decisions. We also reveal model-dependent persona effects, suggesting different architectures implement strategic reasoning through different mechanisms. These findings have practical implications for deployment, as personas alter system behavior in ways that may not be immediately apparent to users, showing that deploying personas requires careful consideration of how identity interacts with payoff optimization in strategic decision making.

### v3

# When Identity Overrides Incentives: Representational Choices as Governance Decisions in Multi-Agent LLM Systems

@@ Abstract.
Large language models are increasingly deployed in multi-agent systems for strategic tasks, yet how design choices such as role-based personas and payoff visibility affect behavior remains poorly understood. We investigate whether LLM agents function as payoff-sensitive strategic actors or as identity-driven role followers. Using a 2x2 factorial experiment (persona presence x payoff visibility) with four models (Qwen-7B/32B, Llama-8B, Mistral-7B), we test 53 environmental policy scenarios in four-agent strategic games. We find that personas suppress payoff-aligned behavior: with personas present, all models achieve near-zero Nash equilibrium in Tragedy-dominant scenarios despite complete payoff information. Nearly every equilibrium reached is Green Transition. Removing personas and providing explicit payoffs are both near-necessary for payoff-aligned behavior, enabling only Qwen models to reach 65–90% equilibrium rates. Our results reveal three behavioral profiles: Qwen adapts to framing, Mistral is disrupted without finding Tragedy equilibrium, and Llama remains near-invariant. We show that the same binary design choice can shift equilibrium attainment by up to 90 percentage points, establishing that representational choices are not implementation details but governance decisions.
@@ 1. Introduction
When a language model is asked to play an industrialist in a multi-agent climate policy game, should it maximize profits or protect the environment? The answer depends not only on the payoff structure of the game, but also on representational choices such as whether the model is given a role-based persona, which can substantially shift behavior.
@@
Multi-agent systems powered by large language models are increasingly being deployed in policy simulation, economic modeling, and negotiation, tasks where agents represent stakeholders with conflicting interests and must coordinate under shared constraints (Merrill and Srivastava, 2025; Li et al., 2024; Bianchi et al., 2024). To make these simulations realistic, agents are typically assigned role-based personas describing their motivations and objectives (Merrill and Srivastava, 2025; Xu et al., 2024b). This design is especially common in environmental and climate related decision-making, where heterogeneous stakeholders interact strategically and coordination failures can lead to persistent social inefficiencies (Hardin, 1968; Ostrom, 1990; Barrett, 2003; Reyes-Cruz et al., 2024). Yet a fundamental question remains unanswered: when agents with role-based identities face scenarios that demand strategic decision making, do they follow explicit payoffs or their assigned roles?
Refer to caption Pipeline diagram showing scenario, representation conditions, multi-agent decisions, and outcome evaluation A left-to-right pipeline diagram illustrating the experimental setup. On the left, a Scenario block describes the situation and defines the game context. Below it, two boxes indicate the manipulated representation conditions: Payoff (Visible or Hidden) and Persona (On or Off). Arrows lead from this setup to four role-assigned agents: Environmentalist, Government, Industrialist, and Citizen. Each agent is shown with a brief reasoning prompt and two binary action choices: Environmentalist chooses CAMPAIGN or NOCAMPAIGN, Government chooses REGULATE or NOREG, Industrialist chooses POLLUTE or CLEAN, and Citizen chooses SUPPORT_GREEN or BUY_CHEAP. The agents’ decisions feed into a central Strategy Profile box representing the joint action. This profile is then evaluated in a final block that determines whether it constitutes a Nash equilibrium, and if so, whether the outcome corresponds to Green Transition or Tragedy of the Commons.
Figure 1. Experimental Pipeline: Each scenario defines a four-agent strategic game with role-specific actions and an underlying payoff structure. Prior to decision-making we vary two factors: whether agents receive role-based persona descriptions and whether payoff information is hidden or explicitly provided. Each agent then produces Chain-of-Thought (CoT) reasoning conditioned on the scenario and its role (illustrated via the speech bubbles) and selects an action. The four agents act simultaneously, producing a joint strategy profile that is evaluated for Nash equilibrium attainment and classified by outcome types (Green Transition or Tragedy of the Commons). CoT reasoning analysis is presented in Appendix I.
The answer matters because it determines what these systems actually produce. A system designed for strategic policy analysis may instead generate role-consistent behavior that ignores encoded incentives (Herr et al., 2024; Duan et al., 2024b; Neumann et al., 2025a). Practitioners relying on these system outputs would have no way to identify this from outside (Bender et al., 2021; Casper et al., 2024). From a governance perspective, these representational choices, such as whether to include personas or reveal payoff information, function as implicit interventions that can shape system-level outcomes. However, the interaction between persona conditioning and payoff visibility in multi-agent strategic settings has not been systematically studied. Existing work has examined personas, payoff visibility, and model differences separately (Neumann et al., 2025a; Duan et al., 2024b; Akata et al., 2025; Huang et al., 2025; Sun et al., 2025; Zhang et al., 2024), but not how they combine to determine behavior.
We investigate this interaction directly. We place four LLM agents (Industrialist, Government, Environmental Activist, and Citizen) in environmental policy scenarios and vary two factors: whether agents receive role-based personas and whether they see explicit payoff tables. Across 53 scenarios, 41 Green-dominant (where coordinated pro-environmental action is the Nash equilibrium (Nash, 1951)) and 12 Tragedy-dominant (where individually beneficial but collectively harmful behavior (Hardin, 1968) is payoff-optimal), we test four models (Qwen-7B/32B, Llama-8B, Mistral-7B) under all four conditions. We also conduct supplementary prompt sensitivity analyses using seven persona formulations per scenario, varying how agent identities are expressed while keeping incentives fixed, to test whether observed behaviors are robust to changes in prompt surface form.
The findings are clear: with personas present, all models achieve near-zero Nash equilibrium rates in Tragedy-dominant scenarios despite complete payoff information, and nearly every equilibrium that is reached corresponds to Green Transition. Removing personas and providing explicit payoffs are both near-necessary to recover payoff-aligned behavior, and even then, only Qwen models respond, reaching 65-90% equilibrium rates. The same intervention that enables equilibrium in Tragedy-dominant scenarios disrupts it in Green-dominant ones: Qwen-32B drops from 96.1% to 16.6% equilibrium attainment when personas are removed and payoffs made visible.
Three distinct behavioral profiles emerge across model families: Qwen adapts to framing, shifting between role-aligned and payoff-aligned behavior; Mistral is strongly affected by representational changes in Green-dominant scenarios but cannot identify the Tragedy equilibrium in any condition; and Llama maintains near-invariant behavior regardless of how the problem is presented.
Our findings establish that the choice of persona design, payoff presentation, and model family jointly determine whether a multi-agent system behaves as a payoff-sensitive strategic actor or an identity-driven role follower. When a language model agent is assigned the role of an industrialist in a climate policy game, whether it maximizes profits or protects the environment is determined not by the game’s incentive structure but by the designer’s choice of prompt.